OpenStudy (anonymous):

what is the sum of 7/10,2/3 and 1/2

OpenStudy (anonymous):

30 13/15

OpenStudy (anonymous):

what do i do for the 2/3

OpenStudy (anonymous):

Make 7/10=21/30, 2/3=20/30, 1/2=15/30, \[21/30+20/30+15/30=56/30=28/15\]